The monthly sales for Yazici batteries were as follows on January 21, February 20, March 16, April 15.May 15,June 18,July 17,August 18, September 22, October 20, November 21, December 24. forcast for the next month (Jan) using the naive method. The forecast for the next period (Jan) using the 3-month moving approach. Using smoothing with a=0.30 and a September forecast of 18.00 the forecast for the next period (Jan) sales.
Answer #1

Answer: Naive method forecast= actual demand of last period

  • Forecast of Jan using naive method= actual demand of december= 24

Answer: Forecast 3 period moving average= average of actual demand of last 3 periods

  • The forecast for the next period (Jan) using the 3-month moving approach= (October 20, November 21, December 24)/3 = 65/3= 21.67

Answer: Forecast simple exponential smoothing, jan= 21.14

  • the formula to be used in Simple Exponential smoothing is

Ft+1= alpha*At + (1-alpha) Ft

At means Actual demand of t'th period, if you want to find out the Forecast through exponential smoothing= forecast of 3rd period = alpha*actual demand of 2nd period +(1-alpha) *forecast demand of 2nd period

remember forecast of September period is 18

Month Demand Exponential forecast, alpha= 0.3
September 22 18.00
October 20 19.20
November 21 19.44
December 24 19.91
January 21.14

Please ask, if you have any doubts through the comment section. Do rate the answer